# Modular Jump Gaussian Processes

### v1

@@ 1 Introduction
We consider processes $Y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ that are smooth for most of the surface except along a manifold of discontinuity which spatially partitions the response $Y(x)$ into two levels depending on $x$. One example of such a process is a computer simulation measuring the average transport time $Y(x)$ of material transport devices given their characteristics $x$ (Park et al.,, 2023). Usually, transport time is low, but when the devices are working at or near capacity transport time suddenly increases – it “jumps”. Other examples of jump dynamics exist in geostatistics (Kim et al.,, 2005), physiology (Vieth,, 1989), ecology (Toms and Lesperance,, 2003), and materials science (Park et al.,, 2022). Our interest is in jumps exhibited by computer simulation experiments, where $Y(x)$ models complex phenomena as a deterministic function of inputs $x$ at potentially great computational expense. In this paper we consider jump processes with only two levels.
@@
When designing computer simulation experiments, one often meta-models a planned experiment of runs of the computer model. That is, gather a collection of $N$ data pairs $(x_{1},y_{1}),\dots,(x_{N},y_{N})$, where $y_{i}=Y(x_{i})$, and obtain a fit that can predict $Y(x)$ at new inputs. Such surrogate models (Santner et al.,, 2003; Gramacy,, 2020) are instrumental to many applications where ready access to runs of the computer model are unavailable or expensive. The canonical surrogate is a Gaussian process (GP; e.g., Sacks et al.,, 1989) because GPs furnish accurate nonlinear predictions with good uncertainty quantification. For this reason they are also popular for many regression applications in machine learning (e.g., Rasmussen and Williams,, 2006) and in geospatial settings (e.g., Banerjee et al.,, 2004). However there are two downsides to GP modeling in many contexts. One is that computational demands for inference scales poorly as $N$ gets large; another is that the typical GP formulation assumes stationarity, in that only relative distances between input locations matter. This means GPs struggle to capture dynamics where modeling fidelity varies in the input space, e.g., where one part is flat and another is wiggly. GPs also struggle if there are jumps of discontinuity in which the level of the response drastically changes.
@@
There is a sizable literature on nonstationary GP modeling. Sauer et al., 2023a provide a recent review. Many such approaches involve imposing hard partitions in the input space (e.g., Kim et al.,, 2005; Pope et al.,, 2021; Payne et al.,, 2020; Gramacy and Lee,, 2008; Luo et al.,, 2023), which could potentially cope with jumps that are axis aligned (like with trees) or are polygonal (Voronoi tesselations). Other approaches involve smoothly warping the inputs into a stationary regime (e.g., Sampson and Guttorp,, 1992; Schmidt and O’Hagan,, 2003; Damianou and Lawrence,, 2013; Sauer et al., 2023b, ), which is often better but would struggle with discontinuity. GPs exist which can handle discontinuities (e.g., Garnett et al.,, 2010; Picheny et al.,, 2019), but they do not assume the response contains distinct levels. Park, (2022) created the “jump GP” (JGP) specifically for data from processes that are piecewise continuous in order to address this gap. JGPs have three main modeling ingredients that are tightly coupled together. The first two are local (GP) modeling (e.g., Gramacy and Apley,, 2015), and latent component (i.e., output level) membership variables. Inference is joint over all unknowns – in particular for the high dimensional latent variable – via modern variants of expectation maximization (EM; Dempster et al.,, 1977), comprising the third ingredient. JGP is a clever, but complex, modeling apparatus that works well out-of-sample when trained on data from response surfaces exhibiting jumps.
@@
Figure 1: All panels show true function $Y(x)$ from Eq. (1) in black. Colored lines represent the mean (solid) and 95% PI (dashed) via (top left) a Global GP, (top right) LAGP, (bottom left) Modular JGP with global GP, and (bottom right) Modular JGP with LAGP. Training data not shown.
@@
Finally, the bottom right panel of Figure 1 combines the feature with local modeling: LAGP applied on $(x_{i},p_{i})$ and output $y_{i}$, but otherwise the same as the feature-expanded GP on the left. We call this the Modular JGP. Except for some elevated uncertainty near the jump points, which is sensible, the fit is nearly perfect: high accuracy, low variance. In the remainder of the paper our goal is to flesh out these ideas, making them more concrete, especially as regards implementation. We begin with review in Section 2. We then describe our main methodological contributions in Sections 3 and 4: optimizing local neighborhoods and cluster-based features. Throughout, we provide illustrations on examples that have the same spirit as Figure 1, but are more realistic. Section 5 provides empirical benchmarking exercises in several variations including on a real simulation from manufacturing. Finally, Section 6 concludes with a brief discussion of limitations and possible future work. All code is provided via Git: https://bitbucket.org/gramacylab/jumpgp.
@@ 2.1 Gaussian Processes
Let $X_{N}$ be an $N\times d$ matrix of input features and $Y_{N}$ be a corresponding $N\times 1$ vector of outputs. Additionally, let $x_{i}^{\top}$ denote the $i^{\mathrm{th}}$ row of $X_{N}$. We consider estimating an unknown function via training data $X_{N}$ and $Y_{N}$ using a GP (Rasmussen and Williams,, 2006; Gramacy,, 2020). A typical GP prior implies that $Y_{N}\sim\mathcal{N}_{N}\bigl{(}\mu,\Sigma(X_{N})\bigr{)}$, where $\mu$ is often assumed to be 0 so that all of the modeling “action” is in $\Sigma(X_{N})$. This setup induces a likelihood function:
@@
The likelihood (2) can be used to learn $\Sigma(X_{N})$ under some restrictions, such as positive definiteness. Often, covariance built from a kernel specifying $\Sigma(X_{N})_{ij}=\Sigma(x_{i},x_{j})$. Kernels that depend only on the distance between inputs yield stationary processes, eliminating all input position information from the model. One example is squared exponential kernel
@@
$\mu_{N}({\mathcal{X}})$ $=\Sigma(\mathcal{X},X_{N})\Sigma(X_{N})^{-1}Y_{N},$ using $\mbox{using }\;\Sigma(\mathcal{X},X_{N})_{ij}$ $=\Sigma(\tilde{x}_{i},x_{j}),$
and $\mbox{and }\quad\Sigma_{N}({\mathcal{X}})$ $=\Sigma(\mathcal{X})-\Sigma(\mathcal{X},X_{N})\Sigma(X_{N})^{-1}\Sigma(\mathcal{X},X_{N})^{\top}.$ (4)
@@
A global GP has two drawbacks: computation becomes prohibitive as $N$ increases, and it does not perform well when training data come from a nonstationary processes. The first can be seen by noting $\Sigma(X_{N})^{-1}$ and $|\Sigma(X_{N})|$ in Eqs. (2) and (4), both $\mathcal{O}(N^{3})$ operations. To see the second, refer back to the discussion of Figure 1 (top-left), illustrating fits on a process with jumps. Consider a single data point at $x=45$, which is part of the level-0 region. The issue with the global GP setup is that two equidistant points, say $x=35$ and $55$, generate the same covariance, even though one point is in the level-0 region and the other is in the level-10 region. However notice that the data is locally stationary, in three spatial regimes, which suggests a divide-and-conquer approach may work well.
@@ 2.2 Local Approximate Gaussian Processes
LAGP (Gramacy and Apley,, 2015) tackles both drawbacks at once: computational expense and stationarity. Observe from Eq. (3) that for two locations $x_{i}$ and $x_{j}$, $\Sigma(x_{i},x_{j})\rightarrow 0$ as $\left\|x_{i}-x_{j}\right\|\rightarrow\infty$. Consider a single new location $x$ where we wish to make a prediction, e.g., a single row of $\mathcal{X}$ in Section 2.1. Note in Eq. (4) that $\Sigma(x,X_{N})$ serves as a weight for $\mu_{N}(\mathcal{X})$; thus, any training $y_{i}$ whose $x_{i}$ is far from $x$ has little influence on $\mu_{N}(\mathcal{X})$. Recognizing this, LAGP chooses a subset of $n\ll N$ “useful” data points $X_{n}(x)\subset X_{N}$ to predict $x$. There are many ways to define “useful”, such as nearest neighbors (NN), mean-squared predictive error, and Fisher information. Throughout we consider LAGP via NN, although it could be worth entertaining other criteria as we discuss briefly in Section 6.
The core idea behind LAGP is that fitting a GP to $X_{n}(x)\subset X_{N}$ with $n\ll N$, along with corresponding $Y_{n}$, is both faster, and more reactive to local dynamics around $x$. We encapsulate that in shorthand as $\text{LAGP}_{n}(x)=\text{GP}\left(x;X_{n}(x),Y_{n}\right)$ after optimizing any hyparameters in $\Sigma(X_{n}(x))$, and $\text{LAGP}_{n}(\mathcal{X})$ for $m$ separate applications covering all of $\mathcal{X}$. The top right panel of Figure 1 shows $\text{LAGP}_{25}(\mathcal{X})$, where the red solid line denotes $\mu_{N}(\mathcal{X})$ and the red dashed line denotes a 95% PI. Since calculations for each ${x}\in\mathcal{X}$ are statistically independent, we don’t get a full covariance structure $\Sigma_{N}(\mathcal{X})$, only a diagonal one with entries $\sigma^{2}_{x}$. The upside, though, is that the calculations are massively parallelizable (Gramacy et al.,, 2014).
LAGP reduces computational costs from $\mathcal{O}(N^{3})$ to $\mathcal{O}(n^{3})$, offering substantial speedups if $n\ll N$. Much work has been done to further decrease the computation time for fixed neighborhood size $n$ (Gramacy and Haaland,, 2016; Sung et al.,, 2018; Cole et al.,, 2021). Besides being faster, LAGP may also be more accurate for data from a nonstationary process. To see this, refer again to Figure 1 (top-right). By excluding points that are far from each $x$, LAGP has captured local trends (at least for the $x<40$ or $x>80$ regions) better than the global GP (top-left). However, LAGP does not perform well in the $8040<x<$ 80 region. This is because these predictive locations have neighbors from across the jumps.
@@ 3 Optimal Neighborhood-size LAGP
The image on the left is formed by evaluating the test function on a dense $101\times 101$ grid in $[0,1]^{2}$. All responses ($y$-values) are deterministic. Observe that there is smooth color variation within the (mostly) red and yellow regions, revealing a strong signal locally (light reds and dark yellows, respectively), even away from the “jumps”. The histogram on the right helps understand the magnitudes of these signals. Notice that red and yellow regimes are have comparable within-color variability and they are well-separated, with no values between 0.4 and 0.6.
@@ 3.1 Searching via validation
Here, we wish to extend that to choosing the best $\hat{n}$. However, we can’t use MSE under a stationary GP for $y$. There are two reasons for this. One is practical: this is not the case we’re interested in – our jump processes are nonstationary. The other is more fundamental. If the process is stationary, then MSE is a decreasing function of $n$: you want as much training data as possible. Since large $n$ is generally intractable, Gramacy and Apley, recommended choosing $n$ as large as computational constraints allow. However, they observed that performance varied with $n$ when the data exhibited telltale signs of nonstationarity, leaving an open question about what to do in practice. So instead of averaging SE, we create a validation exercise which allows us to plug-in an actual $y_{i}$ value from the training data as a proxy.
Let $x_{*}$ be the input in $X_{N}$ that is closest (in terms of Euclidean distance) to $x$. Note that this is also the closest point to $x$ in any $X_{n}(x)$ for $n\geq 1$ and that $x_{*}\equiv X_{1}(x)$. We prefer the simpler $x_{*}$, which is easier on the eyes than $X_{1}(x)$, and allows us to use $y_{*}$ to notate the corresponding element of $Y_{N}$, our proxy for the unknown $y(x)$. Next, let $X_{n}(x_{*})$ denote an $n$-sized neighborhood for predicting at $x_{*}$, i.e., the $n$-nearest points to $x_{*}$ in $X_{N}$, excluding $x_{*}$. Using these quantities, and for any neighborhood size $n$, one may fit $\text{LAGP}_{n}\left(x_{*};X_{n}(x_{*}),Y_{n}\right)$, furnishing a prediction for $y_{*}$ via $\mu_{n}(x_{*})$, streamlining notation. Suppose we do this for each $n$ of interest, and take the best one by squared error:
@@
This $\hat{n}(x_{*})$ is the best neighborhood size for $x_{*}$, which is nearby the location of interest $x$. What remains is to refit the local model at $x$ using the $\hat{n}(x_{*})$-nearest neighbors to $x$, $\mathrm{LAGP}_{\hat{n}(x_{*})}(x;X_{\hat{n}}(x),Y_{\hat{n}})$.
Without care, fitting $(n_{\max}-n_{\min})$-many GPs in pursuit of Eq. (5) could be prohibitively expensive. One option is parallelization, since the calculation for each $n$ is statistically independent, and thus computationally exclusive, of the others. However, we prefer to parallelize instead over the predictive location(s), $x\in\mathcal{X}$, which will be described in Section 5. A computationally thrifty solution takes advantage of the nested structure of neighborhoods as $n$ increases: $X_{n_{\min}}(x_{*})\subseteq X_{n_{\min}+1}(x_{*})\subseteq\dots\subseteq X_{n_{\max}}(x_{*})$. Observe that any two neighborhoods $X_{n}(x_{*})$ and $X_{n-1}(x_{*})$ are identical except for one point: the $n^{\mathrm{th}}$ furthest point from $x_{*}$. Let $x_{n}=X_{n}(x_{*})\setminus X_{n-1}(x_{*})$ denote this new location as the neighborhood increases by one. A visual of these $x_{n}$, selected from $X_{N}$ on a $25\times 25$ grid, is provided in the left panel of Figure 4.
Figure 4: Left: Order of LAGP neighborhood for point $(-0.32,0.17)$; Center: SE by neighborhood size for point $(-0.32,0.17)$; Right: Optimal neighbors LAGP (OLAGP) fit. Blue dots indicate two predictive sites and the surrounding black dots are the respective optimal neighborhoods.
Each number indicates the position of $x_{n}$ for some $n$ relative to $x_{*}\in X_{N}$, the closest point to $x=(-0.32,0.17)$.
Now, proceeding in order for $n=n_{\min},\dots,n_{\max}$, suppose we had already calculated $\mathrm{LAGP}_{n-1}(x_{*})$, beginning initially with $n-1\equiv n_{\min}$ via the description in Section 2.2, yielding $\Sigma_{n-1}^{-1}\equiv\Sigma(X_{n-1}(x_{*}))^{-1}$ and similarly for $\log|\Sigma_{n-1}|$. Incorporating the next point, $x_{n}$, involves adding a new row/column to $\Sigma_{n-1}$ to get $\Sigma_{n}$, which may be calculated without new, cubic cost, matrix decompositions by following the partition inverse equations (Barnett,, 1979). Specifically,
@@
Each update is quadratic in $n$, so that the entire processes require $c\sum_{n}^{n_{\max}}n^{2}\in\mathcal{O}(n_{\max}^{3})$ flops. In other words, the computational expense is no greater than that involved in computing the largest LAGP in the search. Given these quantities, $\mu_{n}(x_{*})$ required as in Eq. (4) is again at worst cubic in $n$.
@@
1:$X_{N},Y_{N},x,n_{\min},n_{\max}$
2:$x_{*}=X_{1}(x)$triangleright nearest neighbor of $x$
3:$\{\Sigma_{n_{\min}}^{-1},\log|\Sigma_{n_{\min}}|\}=\mathrm{chol}(\Sigma(X_{n_{\min}}(x_{*})))$triangleright calculate $\Sigma_{n_{\min}}^{-1}$ and $\log|\Sigma_{n_{\min}}|$ using (3)
4:$\{\tau^{2},\theta\}=\mathrm{MLE}(\Sigma_{n_{\min}}^{-1},\log|\Sigma_{n_{\min}}|,Y_{n_{\min}})$triangleright calculate starting MLE using (2)
5:$\mu_{n_{\min}}(x_{*})=\Sigma(x_{*},X_{n_{\min}}(x_{*}))\Sigma_{n_{\min}}^{-1}Y_{n_{\min}}$triangleright make prediction for $x_{*}$ using (4)
@@
7:     $x_{n}=X_{n}(x_{*})\setminus X_{n-1}(x_{*})$triangleright newest point in neighborhood
8:     $\{\Sigma_{n}^{-1},\log|\Sigma_{n}|\}=\mathrm{update}(\Sigma_{n-1}^{-1},\log|\Sigma_{n-1}|,x_{n})$triangleright calculate new $\Sigma_{n}^{-1}$ and $|\Sigma_{n}|$ using (6)
@@
10:         $\{\tau^{2},\theta\}=\mathrm{MLE}(\Sigma_{n}^{-1},\log|\Sigma_{n}|,Y_{n})$triangleright recalculate MLE using (2)
@@
12:     $\mu_{n}(x_{*})=\Sigma(x_{*},X_{n}(x_{*}))\Sigma_{n}^{-1}Y_{n}$triangleright make prediction for $x_{*}$ using (4)
@@
14:$\hat{n}(x_{*})=\underset{n\in\{n_{\min},\dots,n_{\max}\}}{\mathrm{argmin}}\mathrm{SE}\left(y_{*},\mu_{n}(x_{*})\right)$triangleright choose best neighborhood size using (5)
15:return $\mathrm{LAGP}_{\hat{n}(x_{*})}(x;X_{\hat{n}}(x),Y_{\hat{n}}))$triangleright make prediction for $x$ with $\hat{n}(x_{*})$
@@
Continuing our illustration in Figure 4, the center panel provides a view of the SE-values encountered via Eq. (5) using the setup on the left, for $x=(-0.32,0.17)$, as $n$ ranges from $n_{\min}=12$ to $n_{\max}=160$. The result is $\hat{n}(x)=43$, and the indices for these locations are indicated with white text. Notice how $\hat{n}(x)$ is chosen such that $X_{\hat{n}}(x)$ is as big as possible while residing entirely in the red-response area. This process must be repeated for $x\in\mathcal{X}$ of interest, whereas only one is illustrated in the figure. Some of the details of how we recommend doing this are deferred to Section 5. A for loop around Algorithm 1 is one simple option, and is the basis of the illustration described next.

### v2

@@ 1 Introduction
We consider processes $Y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ that are smooth for most of the surface except along a manifold of discontinuity which spatially [^1] Throughout the paper, we use “spatial” with reference to the input space, not “geospatial” locations on Earth. partitions the response $Y(x)$ into two levels depending on $x$. One example of such a process is a computer simulation of an automated material handling vehicle (AMHV) transport system measuring the average transport time $Y(x)$ of material transport devices given their characteristics $x$ (Park et al.,, 2023). Usually, transport time is low, but when the devices are working at or near capacity transport time suddenly increases – it “jumps”. Other examples of jump dynamics exist in geostatistics (Kim et al.,, 2005), physiology (Vieth,, 1989), ecology (Toms and Lesperance,, 2003), and materials science (Park et al.,, 2022). We narrow our focus to jumps exhibited by computer simulation experiments, where $Y(x)$ models complex phenomena as a deterministic function of inputs $x$ at potentially great computational expense, although we expect our idea to be applicable to any data with jumps in the response. In this paper we consider jump processes with only two levels.
@@
When designing computer simulation experiments, one often meta-models a planned experiment of runs of the computer model. That is, gather a collection of $N$ data pairs $(x_{1},y_{1}),\dots,(x_{N},y_{N})$, where $y_{i}=Y(x_{i})$, and obtain a fit that can predict $Y(x)$ at new inputs. Such surrogate models (Santner et al.,, 2003; Gramacy,, 2020) are instrumental to many applications where ready access to runs of the computer model are unavailable or expensive. The canonical surrogate is a Gaussian process (GP; e.g., Sacks et al.,, 1989) because GPs furnish accurate nonlinear predictions with good uncertainty quantification. For this reason they are also popular for many regression applications in machine learning (e.g., Rasmussen and Williams,, 2006) and geospatial settings (e.g., Banerjee et al.,, 2004), where GPs are known as “kriging” (Schabenberger and Gotway,, 2017; Zimmerman and Ver Hoef,, 2024). There are two downsides to GP modeling in many contexts. One is that computational demands for inference scale poorly as $N$ gets large; another is that the typical GP formulation assumes stationarity, in that only relative distances between input locations matter. This means GPs struggle to capture dynamics where modeling fidelity varies in the input space, e.g., where one part is flat and another is wiggly. GPs also struggle if there are jumps of discontinuity in which the level of the response drastically changes.
@@
There is a sizable literature on nonstationary GP modeling. Sauer et al., 2023a provide a recent review. Many such approaches involve imposing hard partitions in the input space (e.g., Kim et al.,, 2005; Pope et al.,, 2021; Payne et al.,, 2020; Gramacy and Lee,, 2008; Luo et al.,, 2023), which could potentially cope with jumps that are axis aligned (like with trees) or are polygonal (Voronoi tessellations). Other approaches involve smoothly warping the inputs into a stationary regime (e.g., Sampson and Guttorp,, 1992; Schmidt and O’Hagan,, 2003; Damianou and Lawrence,, 2013; Sauer et al., 2023b, ), which is often better but would struggle with discontinuity. GPs exist which can handle discontinuities (e.g., Garnett et al.,, 2010; Picheny et al.,, 2019), but they do not assume the response contains distinct levels. Park, (2022) created the “jump GP” (JGP) specifically for data from processes that are piecewise continuous in order to address this gap. JGPs have three main modeling ingredients that are tightly coupled together. The first two are local (GP) modeling (e.g., Gramacy and Apley,, 2015), and latent component (i.e., output level) membership variables. Inference is joint over all unknowns – in particular for the high dimensional latent variable – via modern variants of expectation maximization (EM; Dempster et al.,, 1977), comprising the third ingredient. JGP is a clever, but complex, modeling apparatus that works well out-of-sample when trained on data from response surfaces exhibiting jumps.
@@
Figure 1: All panels show true function $Y(x)$ from Eq. (1) in black. Colored lines represent the mean (solid) and 95% PI (dashed) via (top left) a global GP, (top right) LAGP, (bottom left) MJGP with global GP, and (bottom right) MJGP with LAGP. Training data not shown.
@@
Finally, the bottom right panel of Figure 1 combines the feature with local modeling: LAGP applied on $(x_{i},p_{i})$ and output $y_{i}$, but otherwise the same as the feature-expanded GP on the left. We call this the Modular JGP (MJGP). Except for some elevated uncertainty near the jump points, which is sensible, the fit is nearly perfect: high accuracy, low variance. In the remainder of the paper our goal is to flesh out these ideas, making them more concrete, especially as regards implementation. We begin with review in Section 2. We then describe our main methodological contributions in Sections 3 and 4: optimizing local neighborhoods and cluster-based features. Throughout, we provide illustrations on examples that have the same spirit as Figure 1, but are more realistic. Section 5 provides empirical benchmarking exercises in several variations including on a real simulation from manufacturing. Finally, Section 6 concludes with a brief discussion of limitations and possible future work. All code is provided via Git: https://bitbucket.org/gramacylab/jumpgp.
@@ 2.1 Gaussian Processes
Let $X_{N}$ be an $N\times d$ matrix of input features and $Y_{N}$ be a corresponding $N\times 1$ vector of outputs. We assume $X_{N}$ has been pre-scaled to the unit $d$-cube. Additionally, let $x_{i}^{\top}$ denote the $i^{\mathrm{th}}$ row of $X_{N}$. We consider estimating an unknown function via training data $X_{N}$ and $Y_{N}$ using a GP (Rasmussen and Williams,, 2006; Gramacy,, 2020). A typical GP prior implies that $Y_{N}\sim\mathcal{N}_{N}\bigl{(}\mu,\Sigma(X_{N})\bigr{)}$, where $\mu$ is often assumed to be 0 so that all of the modeling “action” is in $\Sigma(X_{N})$. This setup induces a likelihood function:
@@
The likelihood (2) can be used to learn $\Sigma(X_{N})$ under some restrictions, such as positive definiteness. Often, covariance is built from a kernel specifying $\Sigma(X_{N})_{ij}=\Sigma(x_{i},x_{j})$. Kernels that depend only on the distance between inputs yield stationary processes, eliminating all input position information from the model. One example is squared exponential kernel
@@
$\mu_{N}({\mathcal{X}})$ $=\Sigma(\mathcal{X},X_{N})\Sigma(X_{N})^{-1}Y_{N},$ $\mbox{using }\;\Sigma(\mathcal{X},X_{N})_{ij}$ $=\Sigma(\tilde{x}_{i},x_{j}),$
$\mbox{and }\quad\Sigma_{N}({\mathcal{X}})$ $=\Sigma(\mathcal{X})-\Sigma(\mathcal{X},X_{N})\Sigma(X_{N})^{-1}\Sigma(\mathcal{X},X_{N})^{\top}.$ (4)
@@
A global GP has two drawbacks: computation becomes prohibitive as $N$ increases, and it does not perform well when training data come from nonstationary processes. The first can be seen by noting $\Sigma(X_{N})^{-1}$ and $|\Sigma(X_{N})|$ in Eqs. (2) and (4), both $\mathcal{O}(N^{3})$ operations. To see the second, refer back to the discussion of Figure 1 (top-left), illustrating fits on a process with jumps. Consider a single data point at $x=45$, which is part of the level-0 region. The issue with the global GP setup is that two equidistant points, say $x=35$ and $55$, generate the same covariance, even though one point is in the level-0 region and the other is in the level-10 region. However notice that the data is locally stationary, in three spatial regimes, which suggests a divide-and-conquer approach may work well.
@@ 2.2 Local Approximate Gaussian Processes
LAGP (Gramacy and Apley,, 2015) tackles both drawbacks at once: computational expense and stationarity. Observe from Eq. (3) that for two locations $x_{i}$ and $x_{j}$, $\Sigma(x_{i},x_{j})\rightarrow 0$ as $\left\|x_{i}-x_{j}\right\|\rightarrow\infty$. Consider a single new location $x$ where we wish to make a prediction, e.g., a single row of $\mathcal{X}$ in Section 2.1. Note in Eq. (4) that $\Sigma(x,X_{N})$ serves as a weight for $\mu_{N}(\mathcal{X})$; thus, any training $y_{i}$ whose $x_{i}$ is far from $x$ has little influence on $\mu_{N}(\mathcal{X})$. Recognizing this, LAGP chooses a subset of $n\ll N$ “useful” data points $X_{n}(x)\subset X_{N}$ to predict $x$. There are many ways to define “useful”, such as nearest neighbors (NN), mean-squared predictive error, and Fisher information. Throughout we consider LAGP via NN, although it could be worth entertaining other criteria as we discuss briefly in Section 6. LAGP is not unique in its use of a small selection nearby points. Other, similar methods include nearest neighbor GPs (Datta et al., 2016a, ; Datta et al., 2016b, ), inducing points (Snelson and Ghahramani,, 2005; Banerjee et al.,, 2008), or kernels and basis functions with compact support (Nychka et al.,, 2015; Kaufman et al.,, 2011). However, LAGP is unique in its focus on points nearby to the predictive location $x$, as opposed to globally or for all $X_{N}$.
The core idea behind LAGP is that fitting a GP to $X_{n}(x)\subset X_{N}$ with $n\ll N$, along with corresponding $Y_{n}$, is both faster, and more reactive to local dynamics around $x$. We encapsulate that in shorthand as $\text{LAGP}_{n}(x)=\text{GP}\left(x;X_{n}(x),Y_{n}\right)$ after optimizing any hyperparameters in $\Sigma(X_{n}(x))$, and $\text{LAGP}_{n}(\mathcal{X})$ for $m$ separate applications covering all of $\mathcal{X}$. The top right panel of Figure 1 shows $\text{LAGP}_{25}(\mathcal{X})$, where the red solid line denotes $\mu_{N}(\mathcal{X})$ and the red dashed line denotes a 95% PI. Since calculations are computationally exclusive for each ${x}\in\mathcal{X}$, we don’t get a full covariance structure $\Sigma_{N}(\mathcal{X})$, only a diagonal one with entries $\sigma^{2}_{x}$. The upside, though, is that the calculations are massively parallelizable (Gramacy et al.,, 2014).
LAGP reduces computational costs from $\mathcal{O}(N^{3})$ to $\mathcal{O}(n^{3})$, offering substantial speedups if $n\ll N$. Much work has been done to further decrease the computation time for fixed neighborhood size $n$ (Gramacy and Haaland,, 2016; Sung et al.,, 2018; Cole et al.,, 2021). Besides being faster, LAGP may also be more accurate for data from a nonstationary process. To see this, refer again to Figure 1 (top-right). By excluding points that are far from each $x$, LAGP has captured local trends (at least for the $x<40$ or $x>80$ regions) better than the global GP (top-left). However, LAGP does not perform well in the $40<x<80$ region. This is because these predictive locations have neighbors from across the jumps.
@@ 3 Optimal Neighborhood-size LAGP
The image on the left is formed by evaluating the test function on a dense $101\times 101$ grid in $0.5,0.5]^{2}$. All responses ($y$-values) are deterministic. Observe that there is smooth color variation within the (mostly) red and yellow regions, revealing a strong signal locally (light reds and dark yellows, respectively), even away from the “jumps”. The histogram on the right helps understand the magnitudes of these signals. Notice that red and yellow regimes are have comparable within-color variability and they are well-separated, with no values between 0.4 and 0.6.
@@ 3.1 Searching via validation
Here, we wish to extend that to choosing the best $\hat{n}$. However, we can’t use MSE under a stationary GP for $y$. There are two reasons for this. One is practical: this is not the case we’re interested in – our jump processes are nonstationary. The other is more fundamental. If the process is stationary, then MSE is a decreasing function of the the subset size $n$. In other words, you want to condition on as much data as possible. Since large $n$ is generally intractable, Gramacy and Apley, recommended choosing $n$ as large as computational constraints allow. However, they observed that performance varied with $n$ when the data exhibited telltale signs of nonstationarity, leaving an open question about what to do in practice. So instead of averaging SE, we create a validation exercise which allows us to plug-in an actual $y_{i}$ value from the training data as a proxy.
Let $x_{*}$ be the input in $X_{N}$ that is closest (in terms of Euclidean distance) to $x$. Note that this is also the closest point to $x$ in any $X_{n}(x)$ for $n\geq 1$ and that $x_{*}\equiv X_{1}(x)$. We prefer the simpler $x_{*}$, which is easier on the eyes than $X_{1}(x)$. Analogously, let $y_{*}$ be the element in $Y_{N}$ corresponding to $x_{*}$. We use $y_{*}$ as a proxy for the value of $y(x)$, the mean response at new location $x$, which is unknown. Next, let $X_{n}(x_{*})$ denote an $n$-sized neighborhood for predicting at $x_{*}$, i.e., the $n$-nearest points to $x_{*}$ in $X_{N}$, excluding $x_{*}$. Using these quantities, and for any neighborhood size $n$, one may fit $\text{LAGP}_{n}\left(x_{*};X_{n}(x_{*}),Y_{n}\right)$, furnishing a prediction for $y_{*}$. Specifically, let $\mu_{n}(x_{*})$ denote the predictive mean of $x_{*}$ calculated with neighborhood size $n$. Suppose we do this for each $n$ of interest, and take the best one by squared error:
@@
[…]
Without care, fitting $(n_{\max}-n_{\min})$-many GPs in pursuit of Eq. (5) could be prohibitively expensive. One option is parallelization, since the calculation for each $n$ is computationally exclusive of the others. However, we prefer to parallelize instead over the predictive location(s), $x\in\mathcal{X}$, which will be described in Section 5. A computationally thrifty solution takes advantage of the nested structure of neighborhoods as $n$ increases: $X_{n_{\min}}(x_{*})\subseteq X_{n_{\min}+1}(x_{*})\subseteq\dots\subseteq X_{n_{\max}}(x_{*})$. Observe that any two neighborhoods $X_{n}(x_{*})$ and $X_{n-1}(x_{*})$ are identical except for one point: the $n^{\mathrm{th}}$ furthest point from $x_{*}$. Let $x_{n}=X_{n}(x_{*})\setminus X_{n-1}(x_{*})$ denote this new location as the neighborhood increases by one. A visual of these $x_{n}$, selected from $X_{N}$ on a $25\times 25$ grid, is provided in the left panel of Figure 4.
Figure 4: Left: Order of LAGP neighborhood for point (0.32,0.17)(-0.32,0.17); Center: SE by neighborhood size for point (0.32,0.17)(-0.32,0.17); Right: Optimal neighbors LAGP (OLAGP) fit. Blue dots indicate two predictive sites and the surrounding black dots are the respective optimal neighborhoods.
Each number indicates the position of $x_{n}$ for some $n$ relative to $x_{*}\in X_{N}$, the closest point to x=(0.32,0.17)x=(-0.32,0.17).
Now, proceeding in order for $n=n_{\min},\dots,n_{\max}$, suppose we had already calculated $\mathrm{LAGP}_{n-1}(x_{*})$, beginning initially with $n-1\equiv n_{\min}$ via the description in Section 2.2, yielding $\Sigma_{n-1}^{-1}\equiv\Sigma(X_{n-1}(x_{*}))^{-1}$ and similarly for $\log|\Sigma_{n-1}|$. Incorporating the next point, $x_{n}$, involves adding a new row/column to $\Sigma_{n-1}$ to get $\Sigma_{n}$, which may be calculated without new cubic cost matrix decompositions by following the partition inverse equations (Barnett,, 1979). Specifically,
@@
Each update is quadratic in $n$, so that the entire process requires $c\sum_{n}^{n_{\max}}n^{2}\in\mathcal{O}(n_{\max}^{3})$ flops. In other words, the computational expense is no greater than that involved in computing the largest LAGP in the search. Given these quantities, $\mu_{n}(x_{*})$ required as in Eq. (4) is again at worst cubic in $n$.
@@
[…]
@@
7:  $x_{n}=X_{n}(x_{*})\setminus X_{n-1}(x_{*})$ $\triangleright$ newest point in neighborhood
8:  $\{\Sigma_{n}^{-1},\log|\Sigma_{n}|\}=\mathrm{update}(\Sigma_{n-1}^{-1},\log|\Sigma_{n-1}|,x_{n})$ $\triangleright$ calculate new $\Sigma_{n}^{-1}$ and $|\Sigma_{n}|$ using (6)
@@
10:   $\{\tau^{2},\theta\}=\mathrm{MLE}(\Sigma_{n}^{-1},\log|\Sigma_{n}|,Y_{n})$ $\triangleright$ recalculate MLE using (2)
@@
12:  $\mu_{n}(x_{*})=\Sigma(x_{*},X_{n}(x_{*}))\Sigma_{n}^{-1}Y_{n}$ $\triangleright$ make prediction for $x_{*}$ using (4)
@@
[…]
@@
Continuing our illustration in Figure 4, the center panel provides a view of the SE-values encountered via Eq. (5) using the setup on the left, for x=(0.32,0.17)x=(-0.32,0.17), as $n$ ranges from $n_{\min}=12$ to $n_{\max}=160$. The result is $\hat{n}(x)=43$, and the indices for these locations are indicated with white text. Notice how $\hat{n}(x)$ is chosen such that $X_{\hat{n}}(x)$ is as big as possible while residing entirely in the red-response area. This process must be repeated for $x\in\mathcal{X}$ of interest, whereas only one is illustrated in the figure. Some of the details of how we recommend doing this are deferred to Section 5. A for loop around Algorithm 1 is one simple option, and is the basis of the illustration described next.